# Generative User-Experience Research for Developing Domain-specific Natural Language Processing Applications

### v1

@@ 5.5 User interface
6 Prototype Evaluation
Our user-driven prototype evaluation aims at assessing prototype accuracy and changes in user value after interacting with the prototype. Unlike concept evaluation at the stage of generative UX research, the evaluation of the prototype is uncontrolled from our side. The combination of the two perspectives (i.e., accuracy and user value) determines the success level of the prototype in the first step toward a system for shared domain knowledge. The evaluation consists of three steps [13]:
  1. 1.
  2. 2.
    experimental: provide the prototype to interact with it through provided tasks for the accuracy assessment and on a free basis;
  3. 3.
    reflection: reuse the same questionnaire and interview scripts to identify how the user experience changes; learn if the system works as expected or surprises them and provides insights.
Table 1 reports a recruited focus group for the prototype evaluation with a diverse set of roles to ensure opinion diversity in the user studies. We invited eighteen study participants from one plant, i.e., domain experts and advanced users of the logging system. Some of the participants were previously recruited for the exploratory generative UX research (see Section 4.2). The users were asked to perform two types of interaction: (1) a near-natural use of the search, i.e., performing a set of tasks to assess the relevance of search results, (2) a natural use, i.e., using the semantic search whenever they have the information need during the execution of their work routines. Fourteen participants who could perform tasks for the relevance assessment were invited to fill in a questionnaire before and after the execution of the task. All participants were encouraged to use the system on demand and were invited to the interviews (of which 8 completed the interviews).
@@ 6 Prototype Evaluation
Total 18 14 8
The section is structured as follows: First, Section 6.1 reports the relevance assessment and evaluation according to the methodology of information retrieval (IR). Second, Section 6.2 summarizes questionnaires and user interviews that measure differences in the user experience before and after interacting with the prototype.
6.1 Relevance assessment
The relevance assessment focuses on the precision-oriented IR evaluation and estimates the relevance of the $N$ top-ranked results. Although such “shallow” evaluation is not complete, it is consistent if users are interested only in top-ranked documents in the domain application [33]. Therefore, the relevance assessment enables a near-natural use of the prototype.
@@ 6.1 Relevance assessment
For the user evaluation, we prepared a list of 100 query topics. We obtained ideas for the query topics during the concept evaluation phase (see Section 4.2) and created the queries based on the records from the plant’s logging system. By following the methodology of [16], we collected diverse queries, phrases of which occurred in the system database at least once. The queries differ by length of query topics, i.e., a query contains 1-4 search terms. Moreover, the queries contain diverse domain terms such as machinery IDs, processes, statuses, and repair steps.
@@
Each of the fourteen participants received 20 queries for evaluation and needed to assess the relevance of the search results on two levels: term and phrase (Figure 13). We distributed a list of query topics among the participants in a way that at least two participants evaluated one query topic. The same queries need to be executed twice to assess the relevance of each of the 20 top-ranked results on the two levels. For a term level, a document is to be labeled as relevant if all query terms occurred in a document as-is or as a synonym independently from their word order or location within the text. For a phrase level, a document is relevant if a query phrase or its meaning occurred in close proximity, i.e., narrow context. For example, for a query topic “problem with a pump,” documents that describe a problem without explicitly mentioning the word “problem,” e.g., “a leakage in a pump,” are to be annotated as relevant.
@@
Figure 13: Study participants need to assess document relevance on two levels: term and phrase match. The difference between the tasks is in the proximity of the relevant terms in a document.
@@
To ensure the validity of the experiment, we calculate inter-rater agreement between the assessors who evaluated document relevance for the same queries. The assessors evaluated approximately half of the distributed queries, i.e., 44 queries for the first task and 16 queries for the second task were assessed at least twice. To measure inter-rater agreement, we calculated Cohen’s kappa for each task [34]. On the term level, Table 2 reports $\kappa_{term}=0.856$, which is a near-perfect agreement between the study participants. On the phrase level, we report $\kappa_{phrase}=0.673$, i.e., the study participants reached a substantial agreement. We assume that lower inter-rater agreement on the phrase-level document relevance depends on two factors: (1) an insufficient number of assessed queries for the second task yielded less reliable kappa computation, (2) assessing phrase-level relevance remains a harder task due to ambiguity in the task description, i.e., what is close proximity. For example, a relevant term of a functional location’s short ID was a record’s attribute, but the remaining relevant terms occurred later in the document. Therefore, a setup could be unclear on how to assess.
@@
Table 2: Inter-rater agreement of Cohen’s kappa for two different tasks: term and phrase level similarity.
@@
To create a test dataset, we combine the assessors’ relevance votes for the query topics assessed more than two times. For each query, we take a list of search results. We sum up votes across assessors and tasks and assign the results as document relevance scores. For example, if two assessors indicate that a document is relevant on both term and phrase levels, then a document gets a relevance score of 4.
@@
We evaluate the prototype and the baselines against the documents collected during the expert assessment of the prototype. We acknowledge that such evaluation may cause certain biases favoring the prototype over baselines. But for the precision evaluation of the prototype, we consider such accuracy approximation sufficient to show how much value the domain users gained by using the prototype. In our evaluation setup of a few user study participants, A/B testing was inappropriate and could have caused statistically irrelevant results [24].
@@
To evaluate the gain in user value after implementing the prototype, we provide two baselines: a keyword search (i.e., exact term match) and BM25. The keyword search is a default SQL-based search from the logging system with which the domain users are familiar. BM25 represents a standard baseline used in IR.
For the experiments, we implement a more flexible version of the keyword search. We lowercase both text to index and search queries, and remove punctuation and stopwords. We retrieve a document if at least one term from a query is found. The documents are ranked by the decreasing number of overlapping terms and then by a timestamp of the records.
Our implementation of BM25 normalizes German text records and accommodates local domain setup. We remove punctuation, and lowercase and lemmatize the text. We retrieve documents if cosine similarity is higher than 0.15, and rank the documents using it and by timestamp.
We implement an optional context expansion in both baselines to evaluate its impact on handling the domain-specific language (see Section 5.2).
@@
We evaluated the accuracy of the prototype at top $N$ records with precision-oriented metrics, i.e., precision $P$@N, and metrics for ranking quality, i.e., mean average precision $mAP$@N [33], mean reciprocal rank $MRR$ [41] and normalized discounted cumulative gain $nDCG$@N [30]. In our experiments, we calculate metrics for the top 5 and 20 records. The combined relevance scores in the test dataset are used in computing $nDCG$@N.
@@ Results
Table 3 reports evaluation results of the proposed prototype and its baselines. Our prototype outperforms all baselines, but even BM25 shows significant performance improvement compared to the original search method in the logging system. Compared to the keyword search, BM25 showed 3-4 times higher values in all metrics, whereas the semantic search improved evaluation by 5-10 times.
Using context expansion improves precision $P$@5 and $P$@20 of both baselines. Figure 14 depicts the influence of context expansion. Interestingly, ranking metrics measured by $MRR$, $mAP$@5, and $mAP$@20, which ignore annotated relevance score of the documents, have slightly dropped. On the contrary, $nDCG$@5 and $nDCG$@20 that take the relevance score into account have increased. It indicates, that context expansion assists not only in retrieving more documents but also the more relevant documents.
@@
6.2 UX evaluation
UX evaluation estimates how user value changes before and after using the prototype, e.g. if users find the information more efficiently and effectively than before. Additionally, through a series of interviews, we collect free-form feedback on using the prototype and ideas for improvements from a user perspective.
@@ 6.2 UX evaluation
The goal of the questionnaires is to compare two “snapshots” of the user experience before and after the relevance assessment, i.e., near-natural use of the prototype. We want to identify the dynamics of how the study participants obtain information by using the prototype compared to old methods of the logging system when only the default search was available. Figure 15 reports the aggregated results of the questionnaires.
Figure 15: Comparison of the user experience in searching for information in the system. The prototype shows a positive trend in both increasing value from the historical data and assisting in the daily work.
For each question in Figure 15, the answers are sorted from the worst to the best option. For all eight questions, we observe a shift in the answers towards the higher-value option after using the prototype. We summarize the improvements made by the prototype as follows. (Q1) Before the participants had access to the prototype, they were rather reading through a list of logs over the past 24-72 hours instead of using the simple search functionality. (Q2-Q3) We observe that the frequency of using search functionality also increased, probably due to the speed-up and better search results of finding answers to their questions. (Q4) The functionality of the context expansion facilitated more efficient search and, consequently, frequent searching for information about them. (Q5) The ultimate goal of the project was to increase the effective lifespan of the logged information (see Figure 7), and we see that after the introduction of the semantic search, the study participants showed more interest in the historic data older than one month. (Q6) We also observed that performing daily operations became more informed due to the increased number of cases when search results helped answer the questions. (Q7) The study participants indicated that to obtain expected results, they needed to rephrase their queries more rarely than before. (Q8) Although we see no almost increase in the complexity of the search queries, the quality of the search results improved (see other questions). We expect the numbers to be more pronounced if operators start using semantic search in a production system and create more semantically complex queries.
@@
User interviews with open-ended questions assist in reflecting on the developed prototype. We collect feedback from interacting with the prototype and ideas for improvements in the next version:
@@
    Did you search with your own search queries? Were there any search queries for which you would have liked to have found more? Was there a search query for which you were particularly impressed by the results or for which you found interesting information?
@@
    Imagine, the prototype is deployed into the productive system. Which functionality would make a difference for you to use the semantic search on a regular basis? What could be improved?
Table 4 reports summaries of the eight interviews and compares a way of finding information by reading records in a list view to using search functionality, i.e., keyword vs. semantic search. Before the prototype evaluation, most of the interviewees preferred using a list view as it was a fast method that complemented well the inefficient old keyword search. Despite low functionality, the old search was still used to find records by keywords. Due to regular interaction with Google search, using any search engine felt natural. After the pilot, on the one hand, a list view was still preferable for specific tasks, i.e., data comparison. On the other hand, semantic search enabled more effective and efficient access to historical data. According to some interviewees, semantic search is a helpful tool to find solutions to similar problems accounted for in the past.
Table 4: Summary of the following questions from the user interviews before and after using the prototype: Which tool do you use to search for information? For which tasks and why? Although the interviewees continue using list view for specific tasks, semantic search becomes a preferred method of finding information of interest.
@@ 7 Discussion
Broad impact
The paper proposes and implements a methodology making generative UX research a part of a pipeline for developing a prototype for NLP application. Our case study shows that learning directly from users or target audience enables identifying their actual problems and needs. Therefore, instead of evaluating hypothetical ideas for an NLP application, the concept generation phase aims at addressing these specific needs in a problem-solving way. The proposed methodology of UX and NLP synergy has a positive impact on developing domain-specific NLP applications. Understanding and evaluating data quality and availability must come in a tight combination with understanding the domain users and their stories.
A timeline depicted in Figure 1 serves as a step-by-step guideline for the generative-driven prototyping of NLP domain applications. The pipeline describes specific tasks at each stage for UX researchers, data scientists, and project managers. Combining the steps in each type of UX research and the questions from the user studies ( Section 4 and Section 6) expands the guideline and equips the UX practitioners with applied examples. Section 4.1, Section 4.2, and Section 5 exemplify data scientists how we performed exploratory data analysis, assessed and solved the challenges for a domain NLP prototype application.
@@
Generative UX research focuses on contextual inquiries aimed at learning from the users how they use an existing product (if any), and which challenges, concerns, and needs they have, thus, exploring opportunities for a new application/feature. To enable immersed domain understanding, contextual inquiries include interviews with the domain users and observations in their environment. Generative research engages domain users from the first step of the development process and targets the identified pain points. This problem-solving methodology increases interest in the final product, sets up already higher user value, and increases user trust in the final application. In our case study, we applied generative UX research to learn how the interviewees use the logging system in their daily operations in the domain of the process industry. The interviews highlighted two problems: (1) the logged text information has an uneven reporting style and suffers from partial incompleteness, (2) users mostly use recently logged information to inefficient access the historical data. The identified problems establish a ground for concept generation.
Exploratory data analysis is essential for ML/NLP applications to understand data availability and quality. In generative research, the preliminary data analysis enables learning about the data constraints that need to be taken into account during the concept generation phase. Moreover, the data analysis facilitates forming hypotheses about patterns of special usage of a domain system, which can be later tested during user interviews. Our exploratory data analysis revealed multiple challenges of the domain text data: (1) semi-structured data records include structured attributes and unstructured text fields, (2) writing style in the records varies from large structured reports to notes quickly taken down with incomplete grammar and syntax, (3) domain language contains rare and domain special terminology, e.g., abbreviations, term shortenings, chemical formulas, numerical values, and codes of the machinery that act as domain terms.
Generated ideas and concepts from UX research help define a scale of a project. Selecting small ideas for a prototype phase speeds up evaluating the credibility of a bigger concept. Based on our UX research, we create an idea of a big-scale system for domain knowledge sharing, e.g., a solution recommendation system that is based on the past experience of plant operation, collected and extracted from the historical records (see Figure 7). For the prototype phase, we selected a small part from this idea and proposed a semantic search to provide effective access to the historical data. The prototype provided us with quick feedback from the domain users and suggested steps for further development.
@@
Prototyping for NLP applications should aim at the UX principle “fail fast, fail often.” When a goal is to evaluate the credibility of a concept in a moderated prototype, the prototype may contain only basic or simplified functionality sufficient for estimating the user value. To handle the domain-specific language in the prototype, we used a combination of two solutions: (1) fastText word embeddings that reliably represent domain compound German terms, and (2) context expansion that adds descriptions to the mentioned machinery IDs. Although these solutions are not from the cutting-edge research of language models and their domain adaption, the prototype enabled reliable evaluation of the user value of the proposed NLP application. Moreover, industrial or domain NLP applications may not contain top-tier research because real-world applications need to deal with more constraints than academic research. Simpler methods do not always mean less effective but more robust and suitable to a specific domain or low-resource language.
Prototype evaluation: Gain in user value
User value should be evaluated thoroughly in domain-specific NLP applications. User value in daily operations in the process industry consists of efficiency, effectiveness, safety, and robustness in the technological processes. An NLP application that assists in operating a plant should improve the quality and completeness of user-generated data to encapsulate user knowledge and experience.
The evaluation methodology employs domain assessors in order to estimate the accuracy of the prototype and its impact on their daily routine. First, we perform a precision-oriented evaluation of the prototype and ask domain users from a focus group to assess the top 20 search results of a given list of queries. The prototype outperformed all baselines, e.g., BM25, by 3-10 times in various metrics. The prototype suits the best for exploring results containing query terms in the original form, shortenings, or synonyms. Retrieving documents based on a phrase-level semantics of a query is a subject for improvement.
Second, we analyze user value with questionnaires and personal interviews. Analysis of the questionnaire difference before and after the prototype evaluation showed a positive trend in improving user experience in working with historical logged data. Figure 15 depicts an increase in such parameters as the speed of acquiring answers to users’ questions and a level of user insights. Moreover, analyses of both questionnaires and interviews indicate that searching becomes a more preferred method of acquiring information than an overview of a list of records. These user studies confirm our hypothesis that effective semantic search facilitates more frequent and effective use of the logged historical data. Figure 7 depicts our second hypothesis that the semantic search increases the completeness and quality of the logged information. Evaluating this hypothesis is not within the scope of this work: improvement in data readability and completeness could be measured only in the long run, as it would require a long-term study using a productive system.

### v4

@@ 4.2.5 User interface
4.3 Evaluation
Our user-driven prototype evaluation aims to assess prototype accuracy and how user utility and effectiveness change after the open interaction with the prototype. We recruited a focus group of seventeen domain experts and advanced users of the logging system from one plant for the prototype evaluation with a diverse set of roles to ensure opinion variance in the user studies (Table 1). Some study participants were previously recruited for the exploratory generative UX research (see Section 4.1.2). The users were asked to perform two types of interaction: (1) perform a set of tasks to assess the relevance of search results, and (2) freely use the system as they see fit during the execution of their work tasks. Fourteen participants who could perform tasks for the relevance assessment were invited to fill in a questionnaire before and after the execution of the task. All participants were encouraged to use the system on demand and were invited to the interviews.
@@ 4.3 Evaluation
Total 17 14 8
The section is structured as follows. Section 4.3.1 reports the relevance assessment and evaluation according to the information retrieval (IR) methodology. Section 4.3.2 summarizes questionnaires and user interviews that measure differences in the user experience before and after interacting with the prototype.
4.3.1 Relevance assessment
The relevance assessment focused on the precision-oriented IR evaluation and estimated the relevance of the $N$ top-ranked results. Although such “shallow” evaluation could be incomplete, it is consistent if users are interested only in top-ranked documents in the domain application [39]. Therefore, the relevance assessment enabled the estimation of the prototype accuracy in daily use.
@@ 4.3.1 Relevance assessment
For the user evaluation, we prepared a list of 100 query topics. We obtained ideas for the query topics during the concept evaluation phase (see Section 4.1.2) and created the queries based on the plant’s logging system records. By following the methodology of [16], we collected diverse queries, phrases of which occurred in the system database at least once. The queries differ by length of query topics, i.e., a query contains 1-4 search terms. Moreover, the queries contain diverse domain terms such as machinery IDs, processes, statuses, and repair steps.
@@
Each of the fourteen participants received 20 queries for evaluation and needed to assess the relevance of the search results on two levels: term and phrase (Figure 13). We distributed a list of query topics among the participants so that at least two participants evaluated one query topic. The same queries needed to be executed twice to assess the relevance of each of the 20 top-ranked results on the two levels. For a term level, a document was to be labeled as relevant if all query terms occurred in a document as-is or as a synonym independently from their word order or location within the text. For a phrase level, a document was relevant if a query phrase or its meaning occurred in close proximity, i.e., in a narrow context. For example, for a query topic “problem with a pump,” the documents that described a problem without explicitly mentioning the word “problem,” e.g., “a leakage in a pump,” were to be annotated as relevant.
@@
Figure 13: Study participants needed to perform a set of tasks in the prototype. The tasks included assessing document relevance on two levels: term and phrase match. The difference between the tasks is in the proximity of the relevant terms in a document.
@@
We calculated inter-rater agreement between the assessors who had evaluated document relevance for the same queries to ensure the experiment’s validity. The assessors evaluated approximately half of the distributed queries, i.e., 44 queries for the first task and 16 queries for the second task were assessed at least twice. We calculated Cohen’s kappa for each task [40] to measure inter-rater agreement. On the term level, Table 2 reports $\kappa_{term}=0.856$, a near-perfect agreement between the study participants. On the phrase level, $\kappa_{phrase}=0.673$, i.e., the study participants reached a substantial agreement. We assume that lower inter-rater agreement on the phrase-level document relevance depends on two factors: (1) an insufficient number of assessed queries for the second task yielded less reliable kappa computation, (2) assessing phrase-level relevance remains a harder task due to ambiguity in the task description, i.e., what is close proximity. For example, a relevant term of a functional location’s code was at the beginning of a record, whereas the remaining relevant terms occurred later in the document. Therefore, a setup could be unclear on how to assess.
@@
Table 2: Inter-rater agreement of Cohen’s kappa for two tasks: term and phrase level similarity. The level of agreement for both tasks is substantial.
@@
To create a test dataset, we combined the assessors’ relevance votes for the query topics assessed more than two times. For each query, we took a list of search results. We summed up votes across assessors and tasks and assigned the results as document relevance scores. For example, if two assessors indicated that a document was relevant on both term and phrase levels, then a document got a relevance score of 4.
@@
We evaluated the prototype and the baselines against the documents collected during the expert assessment of the prototype. We acknowledge that such evaluation may cause certain biases favoring the prototype over baselines. But for the precision evaluation of the prototype, we consider such accuracy approximation sufficient to show how user utility improved by using the prototype. In our evaluation setup of a few user study participants, A/B testing would have been inappropriate and could have caused statistically irrelevant results [25].
@@
After implementing the prototype, we provided two baselines to evaluate how user utility changed: a keyword search (i.e., exact term match) and BM25. The keyword search is a default SQL-based search from the logging system with which the domain users are familiar. BM25 represents a standard baseline used in IR.
We implemented a more flexible version of the keyword search for the experiments. We retrieved a document if at least one term from a query was found. The documents were ranked by the decreasing number of overlapping terms and then by a timestamp of the records.
For BM25, we normalized German text records and accommodated local domain setup with lemmatizing and removing stopwords. We retrieved documents if cosine similarity was higher than 0.15 and ranked them using it and a document’s timestamp.
We implemented an optional context expansion in both baselines to evaluate its impact on handling the domain-specific language (see Section 4.2.2).
@@
We evaluated the accuracy of the prototype at top $N$ records with precision-oriented metrics, i.e., precision $P$@N, and metrics for ranking quality, i.e., mean average precision $mAP$@N [39], mean reciprocal rank $MRR$ [47] and normalized discounted cumulative gain $nDCG$@N [34]. We calculated metrics for the top 5 and 20 records in our experiments. The combined relevance scores in the test dataset were used to compute $nDCG$@N.
@@ Results
Table 3 reports the proposed prototype’s evaluation results and baselines. Our prototype outperformed all baselines, but even BM25 showed significant performance improvement compared to the original search method in the logging system. Compared to the keyword search, BM25 showed 3-4 times higher values in all metrics, whereas the semantic search improved evaluation by 5-10 times.
Using context expansion improved precision $P$@5 and $P$@20 of both baselines. Figure 14 depicts the influence of context expansion. The ranking values measured by $MRR$, $mAP$@5, and $mAP$@20 that ignore the annotated relevance score of the documents have slightly dropped. On the contrary, $nDCG$@5 and $nDCG$@20 that consider the relevance score have increased. The results show that context expansion assisted in retrieving more candidate documents and ranking relevant documents higher.
@@
4.3.2 UX evaluation
UX evaluation estimated how user utility and effectiveness change before and after using the prototype, e.g., if users found the information more efficiently and effectively. Additionally, through interviews, we collected free-form feedback on using the prototype and ideas for improvements from a user perspective.
@@ 4.3.2 UX evaluation
The questionnaires aimed to compare two “snapshots” of the user experience before and after the prototype use. We wanted to see how the study participants obtained information by using the prototype compared to old methods of the logging system when only the default search had been available.
Figure 15: Comparison of the user experience in searching for information in the system. The prototype showed a positive effectiveness and user utility trend using historical data to assist their daily work.
Figure 15 reports the aggregated results of the questionnaires. For each question in Figure 15, the answers are sorted from the worst to the best option. After using the prototype, we observed a shift in the answers towards higher values for all eight questions. We summarized the improvements made by the prototype as follows. (Q1) Before the participants accessed the prototype, they read through a list of logs over the past 24-72 hours instead of using the simple search functionality. (Q2-Q3) We observed that the frequency of search functionality also increased, probably due to the speed-up and better search results of finding answers to their questions. (Q4) The functionality of the context expansion facilitated more efficient search and, consequently, frequent searching for information about them. (Q5) The ultimate goal of the project was to increase the effective lifespan of the logged information (see Figure 7). The chart depicts that after the introduction of the semantic search, the study participants showed more interest in the historical data over one month. (Q6) Performing daily operations became more informed due to the increased number of cases when search results helped answer the questions. (Q7) The study participants indicated they needed to rephrase their queries more rarely than before to obtain expected results. (Q8) Although we saw a marginal increase in the complexity of the search queries, the quality of the search results improved (see other questions). We expect the numbers to be more pronounced if operators use semantic search in a production system and create more semantically complex queries.
@@
User interviews with open-ended questions helped in reflecting on the developed prototype. We collected feedback from interacting with the prototype and ideas for improvements in the next version:
@@
    Did you search with your search queries? Were there any search queries for which you would have liked to have found more? Was there a search query for which you were particularly impressed by the results or for which you found interesting information?
@@
    Imagine the prototype is deployed into the productive system. Which functionality would make a difference in using the semantic search regularly? What could be improved?
Table 4: Summary of the questions from the user interviews before and after using the prototype about the tool preference to search for the information. Although the interviewees continued using list view for specific tasks, semantic search became a preferred method of finding information of interest.
@@
Table 4 reports summaries of the eight interviews and compares finding information by reading records in a list view to using search functionality, i.e., keyword vs. semantic search. Before the prototype evaluation, most interviewees preferred using a list view as it was a fast method that complemented the inefficient old keyword search. Despite low functionality, the old search was still used to find records by keywords. According to some interviewees, using any search engine felt natural due to regular interaction with the Google search engine. After the pilot, on the one hand, a list view was still preferable for specific tasks, i.e., data comparison. On the other hand, semantic search enabled more effective and efficient access to historical data. According to some interviewees, semantic search was a helpful tool to find solutions to similar problems accounted for in the past.
@@ 5 Discussion
The paper reports a case study for which we devised and implemented a methodology of making generative UX research a part of a pipeline for prototyping NLP applications. The case study showed that learning directly from users or target audiences enabled the identification of their problems and needs, which consequently led to effective and efficient prototyping with positive user feedback in the evaluation phase. Therefore, instead of evaluating hypothetical ideas for an NLP application, the concept generation phase addressed these specific needs in a problem-solving way.
We expect the proposed methodology of combining UX and NLP methodologies to impact domain-specific NLP application development positively. Understanding and evaluating data quality and availability must come in a tight combination with understanding the domain users and their stories. A timeline depicted in Figure 1 can serve as a step-by-step guideline for the generative-driven prototyping of NLP domain applications. The pipeline describes specific tasks for UX researchers, data scientists, and project managers at each stage. Combining the steps in each type of UX research and the questions from the user studies (Section 4.1 and Section 4.3) expands the guideline and equips the UX practitioners with applied examples. Section 4.1.1, Section 4.1.2, and Section 4.2 exemplify data scientists how we performed exploratory data analysis, assessed and solved the challenges for a domain NLP prototype application.
We encourage NLP and UX researchers and practitioners to join in evaluating this methodology and reporting their results. Currently, we are testing the methodology in a second case study in which we prototype an NLP application for supporting domain users in logging information. There, we observe and explore how the domain users log information, which steps are hard for them, and which information they enter wrongly or incompletely. Although it is too early to present the results of the case study in full detail, they support our conclusion that using generative UX research for NLP applications strongly improves user acceptance.
@@
Generative UX research focuses on contextual inquiries to learn about the daily routines of domain users and how they use an existing product (if any). Observing and exploring which challenges, concerns, and needs they have enables opportunities for a new NLP application/feature. To facilitate immersed domain understanding, contextual inquiries include interviews with the domain users and observations in their environment. Generative research engages domain users from the first step of the development process and targets the identified pain points. This problem-solving methodology increases interest in the final product and trust in the final application. In our case study, we applied generative UX research to learn how the interviewees use the logging system in their daily operations in the process industry domain. The interviews highlighted two problems: (1) the logged text information has an uneven reporting style and suffers from partial incompleteness, and (2) users mostly used only recently logged information and used inefficient ways to access historical data. The identified problems establish a ground for concept generation.
Exploratory data analysis is essential for ML/NLP applications to understand data availability and quality. In generative research, the preliminary data analysis enables learning about the data constraints that must be considered during the concept generation phase. Moreover, the data analysis facilitates forming hypotheses about patterns of special usage of a domain system, which can be later tested during user interviews. Our exploratory data analysis revealed multiple challenges of the domain text data: (1) semi-structured data records included structured attributes and unstructured text fields, (2) writing style in the records varied from large structured reports to notes quickly taken down with incomplete grammar and syntax, (3) domain language contained rare and domain special terminology, e.g., abbreviations, term shortenings, chemical formulas, numerical values, and codes of the machinery that act as domain terms.
Generated ideas and concepts from UX research help define the scale of a project. Selecting small ideas for a prototype phase speeds up evaluating the credibility of a bigger concept. Based on our UX research, we created an idea for a big-scale system for domain knowledge sharing, e.g., a solution recommendation system that is based on the experience of plant operation, collected and extracted from the historical records (see Figure 7). We selected a small part from this idea for prototyping and proposed a semantic search to provide effective access to the historical data. The prototype provided us with quick feedback from the domain users and suggested steps for further development.
@@
Prototyping for NLP applications should aim at the UX principle “fail fast, fail often.” It means that a primary goal is to evaluate prototype credibility, which leads to a prototype containing only basic or simplified functionality sufficient for estimating the user utility. Moreover, industrial or domain NLP applications may not contain top-tier research because real-world applications need to deal with more constraints than academic research. Simpler methods do not always mean less effective but more robust and suitable to a specific domain or low-resource language. In our case study, we used a combination of two solutions to handle the domain-specific language in the prototype: (1) fastText word embeddings that reliably represent domain compound German terms and (2) context expansion that adds descriptions to the mentioned machinery codes. Although these solutions are not from the cutting-edge research of language models and their domain adaptation, the prototype enabled reliable evaluation of the user utility and usefulness of the proposed NLP application.
Prototype evaluation: effectiveness of the daily task completion
Prototype usefulness and user satisfaction in domain-specific NLP applications are often bound to the personal effectiveness of completing the tasks for which the prototype was intended. Depending on the application, effectiveness can be replaced or extended with efficiency. Since the effectiveness of completing tasks and user utility often depend on the accuracy of an NLP application, the evaluation needs to address a combination of prototype accuracy evaluation when the system is in use and user satisfaction in assisting in their tasks. Quantitative and qualitative analysis of accuracy metrics, questionnaires, and interviews can yield the most complete arguments to decide about the prototype’s success or failure and points for its improvement.
In our case study, user utility in the process industry depends on the effective and efficient acting on occurring problems, which yields safety and quality of the technological processes. The evaluation methodology employs domain assessors to estimate the accuracy of the prototype and its impact on their daily routine. First, we perform a precision-oriented evaluation of the prototype and ask domain users from a focus group to assess the top search results of a given list of queries. Second, we analyze user utility with questionnaires and personal interviews. Analysis of the questionnaire difference before and after the prototype evaluation showed a positive trend in improving user experience working with historical logged data.